# Diffusion Models in Simulation-Based Inference: A Tutorial Review

### v1

@@ 1 Introduction
Simulation-based inference [cranmer2020frontier, SBI;] is concerned with a fundamental question of computational science: given a simulation model of a target system, what can we learn about the unknown parameters $\boldsymbol{\theta}$ of the system given manifest observations $\mathbf{y}$? The same question has been asked from different perspectives, including (Bayesian) statistics [diggle1984monte], inverse problems [tarantola1982inverse], uncertainty quantification [klir2006uncertainty], and deep learning [kingma2014semi]. SBI can be viewed as the intersection of these fields and involves the following three essential components (Figure 1):
@@
Most recently, diffusion models, a generative model family, initially risen to fame for their ability to generate realistic natural images [song2020denoising, HoJai2020a], have permeated the SBI landscape (LABEL:tab:papers). At a surface level, diffusion models can serve as plug-in replacements for any generative model trained on pairs $(\boldsymbol{\theta},\mathbf{y})$. Starting from pure noise, they learn to gradually remove the noise until a realistic target instance (e.g., parameters or observations) remains. Their appeal, however, goes beyond matching existing approaches: they employ a score-based formulation and unconstrained architectures (i.e., any kind of neural network can be used) to provide a level of flexibility that previous model families (e.g., Jacobian-constrained normalizing flows) cannot easily match [chen2025conditional]. This flexibility makes them excel at the often idiosyncratic modeling setups encountered in SBI, where problem formulations may differ in parameterization, conditioning, or training objective, yet ultimately target the same inference goal [geffner2023compositional, sharrock2024sequential, gloeckler2024all].
@@ 2.1 Simulation-Based Inference
SBI is typically motivated as a solution to problems in settings, where all we can do is to simulate the joint model and use the simulations in a smart way to estimate $p(\boldsymbol{\theta}\mid\mathbf{y}_{\text{obs}})$ or $p(\mathbf{y}_{\text{obs}}\mid\boldsymbol{\theta})$. The approach was first introduced by [diggle1984monte], who used kernel density estimation (KDE) for approximating intractable likelihoods. However, it is important to note that SBI is applicable in settings where the likelihood is tractable and where it is not, and hence can be viewed as a general solution to Bayesian computation. In the simplest case, we collect model simulations into a labeled dataset, $\mathcal{D}_{\text{sim}}=\{\mathbf{y}^{(s)},\boldsymbol{\theta}^{(s)}\}_{s=1}^{S}$, along with the unlabeled dataset of real observations, $\mathcal{D}_{\text{obs}}=\{\mathbf{y}_{\text{obs}}^{(r)}\}_{r=1}^{R}$. A key realization in modern SBI was that one can treat the labeled simulations as training data for generative networks, such as diffusion models—the workhorse of this tutorial review. Here, SBI can produce flexible parameter or data generators that can estimate or mimic practically any parametric model. When inferring parameters from real observations or emulating the data model, these generators need to solve a “Sim2Real” problem [elsemuller2025does], which can create extrapolation biases [frazier2024statistical] and calls for a principled Bayesian workflow [li2024amortized].
@@ 2.2 Inference Scenarios and Considerations
Typical inference targets in SBI are the prior [Daras2024-surveyDiffusionModelsInverseProblemsv], likelihood [sharrock2024sequential], posterior [wildberger2023flow], or joint distributions [gloeckler2024all], which can be approximated by diffusion models (LABEL:tab:papers). Each inference target comes with its own trade-offs, depending on the structure of the problem and the dimensionality of $\mathbf{y}$ and $\boldsymbol{\theta}$. Learning the posterior directly offers the most efficient inference, but may require re-training when model components change (e.g., the prior), unless some form of inference-time adaptation is employed [yang2025priorguide]. In contrast, learning the likelihood decouples model training from prior specification. This enables flexible reuse, for instance, in partial pooling or regression models where parameters $\boldsymbol{\theta}$ can be functions of other variables. However, posterior inference with learned likelihoods requires an additional sampling stage and may not be feasible for estimating many posteriors. We discuss learning the joint in Section 4.3.3; for learning the prior in image spaces, we refer to the literature on inverse problems [Daras2024-surveyDiffusionModelsInverseProblemsv].
@@
Model specification $p(\boldsymbol{\theta},\mathbf{y})$ $\boldsymbol{\theta}\sim p(\boldsymbol{\theta}),\quad\mathbf{y}\sim p(\mathbf{y}\mid\boldsymbol{\theta})$ $\boldsymbol{\theta}\sim p(\boldsymbol{\theta}),\quad\mathbf{y}=\texttt{Sim}(\boldsymbol{\theta},\texttt{rng})$
@@
In amortized inference, the diffusion model learns a functional $q(\boldsymbol{\theta}\mid\mathbf{y})$ which—barring any simulation gaps [schmitt2023detecting, frazier2024statistical]—remains valid for any observation $\mathbf{y}_{\text{obs}}$. This approach offloads computation to an upfront training phase and rewards users with instant target estimation over the entire set of observations $\mathcal{D}_{\text{obs}}$. In contrast, direct inference repeats certain computational steps for each observation from scratch. There are two main direct approaches. First, one may learn a likelihood estimator $q(\mathbf{y}\mid\boldsymbol{\theta})$ that can be used in tandem with established Markov chain Monte Carlo (MCMC) [hoffman2014no] samplers for downstream posterior inference. However, even though this likelihood estimator may be trained to remain valid for all pairs $(\boldsymbol{\theta},\mathbf{y})$, posterior inference is not amortized due to the reliance on MCMC for eventually approximating the posterior. Alternatively, one may employ a sequential training scheme that specializes the posterior estimator for a particular observation $\mathbf{y}_{\text{obs}}$ [sharrock2024sequential]. In this case, the approximator learns a tailored function $q(\boldsymbol{\theta}\mid\mathbf{Y}=\mathbf{y}_{\text{obs}})$ that can generate arbitrarily many posterior draws for $\mathbf{y}_{\text{obs}}$ but needs to be re-trained for each new observation.
@@
The simulation and real data budgets are another key consideration in SBI (Table 2). The simulation budget depends on the “computational price” it costs to create a pair $(\boldsymbol{\theta},\mathbf{y})$. In some fields, such as cognitive science, simulations are relatively inexpensive, allowing for a potentially infinite stream of simulations for online training [radev2020bayesflow, fengler2021likelihood]. These fields are well-positioned to benefit from high-throughput SBI [von2022mental]. In other fields, such as fluid dynamics, producing a high-fidelity pair $(\boldsymbol{\theta},\mathbf{y})$ can take multiple days [tumuklu2023temporal], rendering simulation-based training hardly feasible. These domains represent fruitful avenues for future high-impact research, with recent work exploring diffusion models for multi-fidelity parameter estimation [tatsuoka2025multi].
The availability of real data has different constraints than that of simulated data and cannot be increased in general without significant costs. However, the amount of simulated and real data can help decide whether to use an amortized or a non-amortized estimator when many posteriors need to be estimated. Even for a small number of real observations, amortized inference may still be the preferable option, since it serendipitously unlocks powerful diagnostics that require estimating hundreds or thousands of posteriors [talts2018validating, yao2023discriminative, modrak2025simulation, e.g., simulation-based calibration,]. Recent SBI research has also considered semi-supervised approaches that learn from both simulated and real data [huang2023learning, swierc2024domain_npe, elsemuller2025does, mishra2025robust], factoring in the real data budget as a source of training signal.
@@
The dimensionality of the model parameters $\boldsymbol{\theta}$ and the observations $\mathbf{y}$ are further key aspects that, together with the simulation budget, form eight problem archetypes (Table 2). Low-dimensional models typically serve as toy examples for evaluating SBI algorithms (e.g., Gaussian mixtures, two moons). Indeed, the SBI benchmarking suite by [lueckmann2021benchmarking] features ten low-dimensional synthetic toy models (e.g., the maximum dimension of the observation $\mathbf{y}$ is 100). In contrast to synthetic benchmarks, most SBI applications in science come with high-dimensional observations and low-dimensional parameters. This is reflective of the nature of most scientific inquiries: our hypotheses live and die by the sword of some interpretable summaries (e.g., the location and size of an exoplanet) of less interpretable observations (e.g., high-resolution images).
A typical high-dimensional target space arises when image pixels are treated as hidden parameters that need to be estimated from a corrupted image (e.g., Bayesian denoising) or another modality [orozco2025aspire]. These target spaces are the natural habitat of diffusion models, which first rose to prominence as powerful generative models for images [SongErm2020a, HoJai2020a]. Within SBI, however, the focus is not on generating visually pleasing samples but on maximizing the statistical accuracy of the learned distribution [frazier2024statistical]. This emphasis has driven adaptations to model architectures, training objectives, and inference schedules to harmonize with the demands of SBI tasks [geffner2023compositional, wildberger2023flow, holzschuh2024improving, dirmeier2025causal], yielding measurable gains beyond those achievable with image quality–oriented objectives alone.
@@
High Low Low Current SBI benchmarks [lueckmann2021benchmarking]
Low Quijote $N$-body simulations [delaunoy2024low]
High Low High Single-molecule experiments [Dingeldein2025-simulation-basedInferenceSingle-moleculeExperimentsg]
Low Galaxy simulations [zhou2024evaluating]
High High Low 2D Darcy flow [Parno2015-darcy]
Low Groundwater [Hunt2020-groundwater]
High High High Bayesian neural networks [izmailov2021bayesian, BNNs;]
Low Ultrasound imaging [orozco2025aspire]
@@
A final consideration in our non-exhaustive list is the choice of metric for evaluating diffusion models in SBI. Fortunately, since most SBI methods are framed in a Bayesian setting, we can tap into the large pool of general, estimator-agnostic metrics put forward in the Bayesian literature [burkner2023some] and recommended as part of principled Bayesian workflows [gelman2020bayesian]. In general, the Pareto front in SBI moves along two dimensions: efficiency and accuracy. Efficiency can be operationalized via wall-clock time $T$ (or computational complexity) and/or algorithmic complexity necessary for obtaining $L$ samples from one posterior given a budget of $B$ simulations:
@@
For non-amortized (sequential) methods, inference costs $R\cdot T_{\text{inference}}(L)$ grow quickly with increasing number of posteriors $R$ due to the need to re-train the diffusion estimator for each observation, $R\cdot T_{\text{train}}(B)$, and potentially generate new simulations $R\cdot T_{\text{sim}}(B)$ [greenberg2019automatic, griesemer2024active].
@@
where the expectation runs over the true data-generating distribution. This ideal metric is intractable, since the analytic posterior $p(\cdot\mid\mathbf{Y}=\mathbf{y}_{\text{obs}})$ and the full true data-generating distribution $p_{\text{obs}}(\mathbf{y})$ cannot be explicitly evaluated. Thus, proxy metrics that bypass evaluation, such as simulation-based calibration [talts2018validating, yao2023discriminative, modrak2025simulation, SBC;], are typically used and actively researched [sailynoja2025posterior, bansal2025surprising, e.g.,].
Finally, one may argue that posterior predictive checks (PPCs) and metrics such as the Expected Log Predictive Density [piironen2017comparison, ELPD;] may also be used to evaluate diffusion models for SBI. This is certainly valid in silico, where the Bayesian posterior guarantees optimality in posterior predictions [aitchison1975goodness]. However, the relationship between metrics in $\boldsymbol{\theta}$-space and $\mathbf{y}$-space is not always straightforward [scholz2025prediction]. Further, data-space checks with respect to $\mathbf{y}_{\text{obs}}$ inevitably conflate two sources of error: (1) errors due to an inaccurate posterior approximation (e.g., an underexpressive diffusion estimator), and (2) errors due to simulation gaps. Accordingly, good predictive performance in the open world can be achieved by deviating from the target posterior under the potentially misspecified model [lai2024predictive]. Thus, even though PPCs are vital in practical Bayesian workflows [gelman2020bayesian], they are ill-suited as a proxy for measuring posterior accuracy unless they are performed in silico or recast as generalized Bayesian inference by using proper scoring rules or divergence measures in place of the likelihood [gao2023generalized, pacchiardi2024generalized].
@@ 3 Diffusion Models for Simulation-Based Inference
Diffusion models treat target generation as a denoising process [SongErm2020a, HoJai2020a]: Starting from pure noise (usually from a Gaussian), they gradually remove the noise until a realistic target instance remains. Diffusion—the gradual addition of noise until the targets are no longer recognizable—is the opposite process and plays a crucial role in the training of the denoising algorithm. Intuitively, diffusion models record a movie of the target dissolving into noise and then replay it in reverse to recreate the target. Starting with a real target instance $\mathbf{z}_{0}$, the “noising” process at time point $t$ (with initial time $t=0$ and final time $t=1$) performs iterations of the form
@@
Here, the scalar $f_{t}\leq 0$ is a shrinkage term that optionally scales the target towards 0, and $g_{t}$ is the scalar noise standard deviation at time $t$. By making time steps infinitely small, $\Delta t\rightarrow 0$, (3) can be reformulated as a stochastic differential equation (SDE) [SongSoh2020]
@@
For ease of notation, we will also use $t$ for the reverse SDE. The SDE is usually solved numerically in a “one step towards $\mathbf{z}_{0}$, small step away” manner: The first term on the RHS reduces the noise, but the second term adds some fresh noise in every iteration. This counterintuitive strategy often prevents the reverse process from getting stuck in a bad local optimum for $\mathbf{z}_{t}$. The new drift term can also be expressed analytically [anderson1982reverse]:
@@
This ODE is derived from the Fokker-Planck equation, where the factor $1/2$ arises from converting the variance of the Wiener process into a deterministic drift [holderrieth2025flow, for a detailed derivation, see]. Since the ODE also depends on the score, the learning problems for (6) and (8) are closely related.
@@
Beyond denoising score matching, some approaches attempt to estimate the score at the target $\mathbf{z}_{0}$ directly, without dependence on time. This typically requires computing the Jacobian trace of the score, but it enables direct regularization of the score itself to ease training [osada2024local], particularly in likelihood score estimation, where structural properties such as additivity, curvature, and mean-zero behavior are known [jiang2025simulationbasedinferencelangevindynamics]. In regions of low target density, however, score matching may lack sufficient signal to estimate score functions reliably [SongErm2020a].
A popular alternative is flow matching [lipman2023flow, liu2023flow]. It trains a neural network $\hat{u}(\mathbf{z}_{t},t)\approx v(\mathbf{z}_{t},t)$ to approximate the vector field in the probability flow (8). In combination with the noise schedule $\alpha_{t}=1-t$ and $\sigma_{t}=t$ (with end time $T=1$), this leads to the flow-matching loss:
@@
In SBI, we typically want to control the generative process by conditioning on auxiliary information $\mathbf{x}$, such as observations in posterior estimation or parameters in likelihood estimation (i.e., surrogate modeling). The conditional objective can also be related back to the unconditional score [vincent2011connection, LiYua2024], and hence the corresponding objective is a conditional extension of the denoising score matching loss (9):
@@
This equality is true as long as $\log p(\mathbf{z}_{t}\mid\mathbf{z}_{0},\mathbf{x})$ is differentiable with respect to $\mathbf{z}_{t}$ and does not depend on the particular form of $p(\mathbf{z}_{t}\mid\mathbf{z}_{0})$ and $p(\mathbf{z}_{t}\mid\mathbf{z}_{0},\mathbf{x})$ [vincent2011connection]. Intuitively, the score at time $t$ is determined by the target $\mathbf{z}_{0}$ alone, but the conditioning on $\mathbf{x}$ during the reverse process helps steer denoising toward the subsets of the distribution that match the observations or parameters of interest. Thus, conditioning restricts the space of plausible $\mathbf{z}_{0}$: without $\mathbf{x}$, the network must cover all possible clean signals, but with $\mathbf{x}$ it only needs to explain those consistent with the condition.
@@
Given a trained score model $\hat{s}(\mathbf{z}_{t},\mathbf{x},t)\approx\nabla_{\mathbf{z}_{t}}\!\log p(\mathbf{z}_{t}\mid\mathbf{z}_{0},\mathbf{x})$, we can sample from the target distribution $p(\mathbf{z}\mid\mathbf{x})$ by integrating either the reverse SDE (6) or the reverse ODE (8). This continuous formulation allows us to use state-of-the-art SDE and ODE solvers [SongSoh2020]. A typical procedure starts with a draw from the base distribution $\mathbf{z}_{1}\sim p_{1}(\mathbf{z}_{1})$, which is usually a diagonal Gaussian distribution, and then solves the flow from $t=1$ down to $t=0$, producing a draw from the target distribution $\mathbf{z}_{0}\sim p(\mathbf{z}\mid\mathbf{x})$:
@@
where $\Phi$ denotes the flow induced by either the reverse SDE (6) or the probability-flow ODE (8). The need to repeatedly evaluate the neural network within a numerical solver for each random draw is what makes inference with diffusion models slower than single-step methods [papamakarios2021normalizing, e.g., normalizing flows,]. This motivates the use of distillation techniques (Section 5.6.2), which train a separate model to mimic the multi-step diffusion process in a single forward pass, and have so far remained markedly underutilized in SBI.
In addition to sampling, we can evaluate (log) densities $p(\mathbf{z}\mid\mathbf{x})$ by integrating the probability flow ODE (8) from the target back to the base distribution and applying the instantaneous change of variables formula [chen2018neural]:
@@
where the divergence of the velocity field, $\operatorname{div}v(\mathbf{z}_{t},\mathbf{x},t)=\operatorname{Tr}(\nabla_{\mathbf{z}_{t}}v(\mathbf{z}_{t},\mathbf{x},t))$, quantifies how much the flow expands or contracts at each time step. In order to compute the divergence efficiently, we can use an unbiased estimator known as the Skilling-Hutchinson trace estimator [skilling1989eigenvalues, hutchinson1989stochastic]

### v2

@@ 1 Introduction
Simulation-based inference [31, SBI;] is concerned with a fundamental question of computational science: given a simulation model of a target system, what can we learn about the unknown parameters $\boldsymbol{\theta}$ of the system given manifest observations $\mathbf{y}$? The same question has been asked from different perspectives, including (Bayesian) statistics [38], inverse problems [172], uncertainty quantification [95], and deep learning [93]. SBI can be viewed as the intersection of these fields and involves the following three essential components (Figure 1):
@@
Most recently, diffusion models, a generative model family, initially risen to fame for their ability to generate realistic natural images [160, 71], have permeated the SBI landscape (LABEL:tab:papers). At a surface level, diffusion models can serve as plug-in replacements for any generative model trained on pairs $(\boldsymbol{\theta},\mathbf{y})$. Starting from pure noise, they learn to gradually remove the noise until a realistic target instance (e.g., parameters or observations) remains. Their appeal, however, goes beyond matching existing approaches: they employ a score-based formulation and unconstrained architectures (i.e., any kind of neural network can be used) to provide a level of flexibility that previous model families (e.g., Jacobian-constrained normalizing flows) cannot easily match [23]. This flexibility makes them excel at the often idiosyncratic modeling setups encountered in SBI, where problem formulations may differ in parameterization, conditioning, or training objective, yet ultimately target the same inference goal [53, 153, 58].
@@ 2.1 Simulation-Based Inference
SBI is typically motivated as a solution to problems in settings, where all we can do is to simulate the joint model and use the simulations in a smart way to estimate $p(\boldsymbol{\theta}\mid\mathbf{y}_{\text{obs}})$ or $p(\mathbf{y}_{\text{obs}}\mid\boldsymbol{\theta})$. The approach was first introduced by [38], who used kernel density estimation (KDE) for approximating intractable likelihoods. However, it is important to note that SBI is applicable in settings where the likelihood is tractable and where it is not, and hence can be viewed as a general solution to Bayesian computation. In the simplest case, we collect model simulations into a labeled dataset, $\mathcal{D}_{\text{sim}}=\{\mathbf{y}^{(s)},\boldsymbol{\theta}^{(s)}\}_{s=1}^{S}$, along with the unlabeled dataset of real observations, $\mathcal{D}_{\text{obs}}=\{\mathbf{y}_{\text{obs}}^{(r)}\}_{r=1}^{R}$. A key realization in modern SBI was that one can treat the labeled simulations as training data for generative networks, such as diffusion models—the workhorse of this tutorial review. Here, SBI can produce flexible parameter or data generators that can estimate or mimic practically any parametric model. When inferring parameters from real observations or emulating the data model, these generators need to solve a “Sim2Real” problem [43], which can create extrapolation biases [49] and calls for a principled Bayesian workflow [102].
@@ 2.2 Inference Scenarios and Considerations
Typical inference targets in SBI are the prior [33], likelihood [153], posterior [190], or joint distributions [58], which can be approximated by diffusion models (LABEL:tab:papers). Each inference target comes with its own trade-offs, depending on the structure of the problem and the dimensionality of $\mathbf{y}$ and $\boldsymbol{\theta}$. Learning the posterior directly offers the most efficient inference, but may require re-training when model components change (e.g., the prior), unless some form of inference-time adaptation is employed [192]. In contrast, learning the likelihood decouples model training from prior specification. This enables flexible reuse, for instance, in partial pooling or regression models where parameters $\boldsymbol{\theta}$ can be functions of other variables. However, posterior inference with learned likelihoods requires an additional sampling stage and may not be feasible for estimating many posteriors. We discuss learning the joint in Section 4.3.3; for learning the prior in image spaces, we refer to the literature on inverse problems [33].
@@
Joint model $p(\boldsymbol{\theta},\mathbf{y})$ $\boldsymbol{\theta}\sim p(\boldsymbol{\theta}),\quad\mathbf{y}\sim p(\mathbf{y}\mid\boldsymbol{\theta})$ $\boldsymbol{\theta}\sim p(\boldsymbol{\theta}),\quad\mathbf{y}=\texttt{Sim}(\boldsymbol{\theta},\texttt{rng})$
@@
In amortized inference, the diffusion model learns a functional $q(\boldsymbol{\theta}\mid\mathbf{y})$ which—barring any simulation gaps [150, 49]—remains valid for any observation $\mathbf{y}_{\text{obs}}$. This approach offloads computation to an upfront training phase and rewards users with instant target estimation over the entire set of observations $\mathcal{D}_{\text{obs}}$. In contrast, direct inference repeats certain computational steps for each observation from scratch. There are two main direct approaches. First, one may learn a likelihood estimator $q(\mathbf{y}\mid\boldsymbol{\theta})$ that can be used in tandem with established Markov chain Monte Carlo (MCMC) [77] samplers for downstream posterior inference. However, even though this likelihood estimator may be trained to remain valid for all pairs $(\boldsymbol{\theta},\mathbf{y})$, posterior inference is not amortized due to the reliance on MCMC for eventually approximating the posterior. Alternatively, one may employ a sequential training scheme that specializes the posterior estimator for a particular observation $\mathbf{y}_{\text{obs}}$ [153]. In this case, the approximator learns a tailored function $q(\boldsymbol{\theta}\mid\mathbf{Y}=\mathbf{y}_{\text{obs}})$ that can generate arbitrarily many posterior draws for $\mathbf{y}_{\text{obs}}$ but needs to be re-trained for each new observation.
@@
The simulation and real data budgets are another key consideration in SBI (Table 2). The simulation budget depends on the “computational price” it costs to create a pair $(\boldsymbol{\theta},\mathbf{y})$. In some fields, such as cognitive science, simulations are relatively inexpensive, allowing for a potentially infinite stream of simulations for online training [139, 46]. These fields are well-positioned to benefit from high-throughput SBI [183]. In other fields, such as fluid dynamics, producing a high-fidelity pair $(\boldsymbol{\theta},\mathbf{y})$ can take multiple days [179], rendering simulation-based training hardly feasible. These domains represent fruitful avenues for future high-impact research, with recent work exploring diffusion models for multi-fidelity parameter estimation [173].
The availability of real data has different constraints than that of simulated data and cannot be increased in general without significant costs. However, the amount of simulated and real data can help decide whether to use an amortized or a non-amortized estimator when many posteriors need to be estimated. Even for a small number of real observations, amortized inference may still be the preferable option, since it serendipitously unlocks powerful diagnostics that require estimating hundreds or thousands of posteriors [169, 193, 118, e.g., simulation-based calibration,]. Recent SBI research has also considered semi-supervised approaches that learn from both simulated and real data [80, 168, 43, 116], factoring in the real data budget as a source of training signal.
@@
The dimensionality of the model parameters $\boldsymbol{\theta}$ and the observations $\mathbf{y}$ are further key aspects that, together with the simulation budget, form eight problem archetypes (Table 2). Low-dimensional models typically serve as toy examples for evaluating SBI algorithms (e.g., Gaussian mixtures, two moons). Indeed, the SBI benchmarking suite by [114] features ten low-dimensional synthetic toy models (e.g., the maximum dimension of the observation $\mathbf{y}$ is 100). In contrast to synthetic benchmarks, most SBI applications in science come with high-dimensional observations and low-dimensional parameters. This is reflective of the nature of most scientific inquiries: our hypotheses live and die by the sword of some interpretable summaries (e.g., the location and size of an exoplanet) of less interpretable observations (e.g., high-resolution images).
A typical high-dimensional target space arises when image pixels are treated as hidden parameters that need to be estimated from a corrupted image (e.g., Bayesian denoising) or another modality [126]. These target spaces are the natural habitat of diffusion models, which first rose to prominence as powerful generative models for images [164, 71]. Within SBI, however, the focus is not on generating visually pleasing samples but on maximizing the statistical accuracy of the learned distribution [49]. This emphasis has driven adaptations to model architectures, training objectives, and inference schedules to harmonize with the demands of SBI tasks [53, 190, 76, 40], yielding measurable gains beyond those achievable with image quality–oriented objectives alone.
@@
High Low Low Current SBI benchmarks [114]
Low Quijote $N$-body simulations [36]
High Low High Single-molecule experiments [39]
Low Galaxy simulations [205]
High High Low 2D Darcy flow [133]
Low Groundwater [81]
High High High Bayesian neural networks [83, BNNs;]
Low Ultrasound imaging [126]
@@
A final consideration in our non-exhaustive list is the choice of metric for evaluating diffusion models in SBI. Fortunately, since most SBI methods are framed in a Bayesian setting, we can tap into the large pool of general, estimator-agnostic metrics put forward in the Bayesian literature [18] and recommended as part of principled Bayesian workflows [57]. In general, the Pareto front in SBI moves along two dimensions: efficiency and accuracy. Efficiency can be operationalized via wall-clock time $T$ (or computational complexity) and/or algorithmic complexity necessary for obtaining $L$ samples from one posterior given a budget of $B$ simulations:
@@
For non-amortized (sequential) methods, inference costs $R\cdot T_{\text{inference}}(L)$ grow quickly with increasing number of posteriors $R$ due to the need to re-train the diffusion estimator for each observation, $R\cdot T_{\text{train}}(B)$, and potentially generate new simulations $R\cdot T_{\text{sim}}(B)$ [60, 62].
@@
where the expectation runs over the true data-generating distribution. This ideal metric is intractable, since the analytic posterior $p(\cdot\mid\mathbf{Y}=\mathbf{y}_{\text{obs}})$ and the full true data-generating distribution $p_{\text{obs}}(\mathbf{y})$ cannot be explicitly evaluated. Thus, proxy metrics that bypass evaluation, such as simulation-based calibration [169, 193, 118, SBC;], are typically used and actively researched [146, 10, e.g.,].
Finally, one may argue that posterior predictive checks (PPCs) and metrics such as the Expected Log Predictive Density [136, ELPD;] may also be used to evaluate diffusion models for SBI. This is certainly valid in silico, where the Bayesian posterior guarantees optimality in posterior predictions [1]. However, the relationship between metrics in $\boldsymbol{\theta}$-space and $\mathbf{y}$-space is not always straightforward [152]. Further, data-space checks with respect to $\mathbf{y}_{\text{obs}}$ inevitably conflate two sources of error: (1) errors due to an inaccurate posterior approximation (e.g., an underexpressive diffusion estimator), and (2) errors due to simulation gaps. Accordingly, good predictive performance in the open world can be achieved by deviating from the target posterior under the potentially misspecified model [98]. Thus, even though PPCs are vital in practical Bayesian workflows [57], they are ill-suited as a proxy for measuring posterior accuracy unless they are performed in silico or recast as generalized Bayesian inference by using proper scoring rules or divergence measures in place of the likelihood [51, 130].
@@ 3 Diffusion Models for Simulation-Based Inference
Diffusion models treat target generation as a denoising process [164, 71]: Starting from pure noise (usually from a Gaussian), they gradually remove the noise until a realistic target instance remains. Diffusion—the gradual addition of noise until the targets are no longer recognizable—is the opposite process and plays a crucial role in the training of the denoising algorithm. Intuitively, diffusion models record a movie of the target dissolving into noise and then replay it in reverse to recreate the target. Starting with a real target instance $\mathbf{z}_{0}$, the “noising” process at time point $t$ (with initial time $t=0$ and final time $t=1$) performs iterations of the form
@@
Here, the scalar $f_{t}\leq 0$ is a shrinkage term that optionally scales the target towards 0, and $g_{t}$ is the scalar noise standard deviation at time $t$. By making time steps infinitely small, $\Delta t\rightarrow 0$, (3) can be reformulated as a stochastic differential equation (SDE) [166]
@@
For ease of notation, we will also use $t$ for the reverse SDE. The SDE is usually solved numerically in a “one step towards $\mathbf{z}_{0}$, small step away” manner: The first term on the RHS reduces the noise, but the second term adds some fresh noise in every iteration. This counterintuitive strategy often prevents the reverse process from getting stuck in a bad local optimum for $\mathbf{z}_{t}$. The new drift term can also be expressed analytically [2]:
@@
This ODE is derived from the Fokker-Planck equation, where the factor $1/2$ arises from converting the variance of the Wiener process into a deterministic drift [74, for a detailed derivation, see]. Since the ODE also depends on the score, the learning problems for (6) and (8) are closely related.
@@
Beyond denoising score matching, some approaches attempt to estimate the score at the target $\mathbf{z}_{0}$ directly, without dependence on time. This typically requires computing the Jacobian trace of the score, but it enables direct regularization of the score itself to ease training [129], particularly in likelihood score estimation, where structural properties such as additivity, curvature, and mean-zero behavior are known [86]. In regions of low target density, however, score matching may lack sufficient signal to estimate score functions reliably [164].
A popular alternative is flow matching [106, 107]. It trains a neural network $\hat{u}(\mathbf{z}_{t},t)\approx v(\mathbf{z}_{t},t)$ to approximate the vector field in the probability flow (8). In combination with the noise schedule $\alpha_{t}=1-t$ and $\sigma_{t}=t$ (with end time $T=1$), this leads to the flow-matching loss:
@@
In SBI, we typically want to control the generative process by conditioning on auxiliary information $\mathbf{x}$, such as observations in posterior estimation or parameters in likelihood estimation (i.e., surrogate modeling). The conditional objective can also be related back to the unconditional score [181, 103], and hence the corresponding objective is a conditional extension of the denoising score matching loss (9):
@@
This equality is true as long as $\log p(\mathbf{z}_{t}\mid\mathbf{z}_{0},\mathbf{x})$ is differentiable with respect to $\mathbf{z}_{t}$ and does not depend on the particular form of $p(\mathbf{z}_{t}\mid\mathbf{z}_{0})$ and $p(\mathbf{z}_{t}\mid\mathbf{z}_{0},\mathbf{x})$ [181]. Intuitively, the score at time $t$ is determined by the target $\mathbf{z}_{0}$ alone, but the conditioning on $\mathbf{x}$ during the reverse process helps steer denoising toward the subsets of the distribution that match the observations or parameters of interest. Thus, conditioning restricts the space of plausible $\mathbf{z}_{0}$: without $\mathbf{x}$, the network must cover all possible clean signals, but with $\mathbf{x}$ it only needs to explain those consistent with the condition.
@@
Given a trained score model $\hat{s}(\mathbf{z}_{t},\mathbf{x},t)\approx\nabla_{\mathbf{z}_{t}}\!\log p(\mathbf{z}_{t}\mid\mathbf{z}_{0},\mathbf{x})$, we can sample from the target distribution $p(\mathbf{z}\mid\mathbf{x})$ by integrating either the reverse SDE (6) or the reverse ODE (8). This continuous formulation allows us to use state-of-the-art SDE and ODE solvers [166]. A typical procedure starts with a draw from the base distribution $\mathbf{z}_{1}\sim p_{1}(\mathbf{z}_{1})$, which is usually a diagonal Gaussian distribution, and then solves the flow from $t=1$ down to $t=0$, producing a draw from the target distribution $\mathbf{z}_{0}\sim p(\mathbf{z}\mid\mathbf{x})$:
@@
where $\Phi$ denotes the flow induced by either the reverse SDE (6) or the probability-flow ODE (8). The need to repeatedly evaluate the neural network within a numerical solver for each random draw is what makes inference with diffusion models slower than single-step methods [132, e.g., normalizing flows,]. This motivates the use of distillation techniques (Section 5.6.2), which train a separate model to mimic the multi-step diffusion process in a single forward pass, and have so far remained markedly underutilized in SBI.
In addition to sampling, we can evaluate (log) densities $p(\mathbf{z}\mid\mathbf{x})$ by integrating the probability flow ODE (8) from the target back to the base distribution and applying the instantaneous change of variables formula [22]:
@@
where the divergence of the velocity field, $\operatorname{div}v(\mathbf{z}_{t},\mathbf{x},t)=\operatorname{Tr}(\nabla_{\mathbf{z}_{t}}v(\mathbf{z}_{t},\mathbf{x},t))$, quantifies how much the flow expands or contracts at each time step. In order to compute the divergence efficiently, we can use an unbiased estimator known as the Skilling-Hutchinson trace estimator [159, 82]